How do you find x and y intercepts of a logarithmic function?
Ex. y=log(-2x+2)-1

Found 2 solutions by FrankM, mananth:
Answer by FrankM(1040)

y=log(-2x+2)-1
For the y intercept, X=0, and
y=log(-2(0)+2)-1
y=log(2)-1
The log of 2 is .301, and .301 - 1 = -.699
To solve for y=0
0=log(-2x+2)-1
1=log(-2x+2)
rearrange
10^(-2x+2)=1
Any non-zero number raised to the zero power equals 1, so
(-2x+2)=0
and X=-4
Here is the Very Cool Graph(tm) for this function


Answer by mananth(16946)

y=log(-2x+2)-1
For y intercept plug x=0
y = log2-1
y intercept = -0.69
For x intercept plug y =0

log(-2x+2)-1=0
log(-2x+2)=1
-2x+2 = 10
-2x = 8
x=-4
x intercept = -4